# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ 6 Conclusion
In this work, we connect multi-accurate learning and show how off-the-shelf multi-accurate learning can be used for conditional average treatment effect estimation that is robust to unknown covariate shift. Although we empirically compare to more “state of the art” causal machine learning, these methods were designed for different purposes. Important directions for future work include “best-of-both-worlds” guarantees on both robustness and efficiency by improving variance reduction properties of Multi-CATE. A finer-grained analysis of the statistical implications of algorithmic implementations of boosting could also be relevant, in addition to improving hyperparameter tuning in the causal setting. In our work, we focus on establishing robustness properties.
@@ Appendix A Notation summary
$\mathcal{D}{\text{obs}}$ Observational dataset
$\mathcal{D}{\text{rct}}$ Randomized controlled trial dataset
@@ Appendix B Details on algorithms
elsereturn $\tilde{p}=p_{k}$// return when miscalibration small
@@ Appendix C Proofs
(2) Suppose any of $\hat{\mu}_{1},\hat{\mu}_{0}$ are not consistent estimators and $e_{1}^{-1},(1-e_{1})^{-1}\in\mathcal{H}.$ The implications of multi-accuracy with respect to $\mathcal{H}$ relate to the classical doubly-robust estimator:
@@ D.1 Data and Software
In all experiments, (initial) T-learner and DR-learner are post-processed using the MCBoost algorithm as implemented in the mcboost (0.4.2) [Pfisterer et al., 2021] package. More concretely, we make use of boosting for degree-2 multi-calibration, a (slightly) stronger notion than multi-accuracy, but computationally less demanding than full multi-calibration Gopalan et al. [2022b]. The hyperparameter settings used for post-processing are listed as part of the following detailed presentation of the experiments (Table 2b and 12b).
@@ D.2.1 Setup
    and used to simulate externally shifted observational data $\mathcal{D}_{os-shift}$ or shifted randomized control trial (RCT) data, $\mathcal{D}_{rct}$ (where $e(x)=0.5$), depending on the simulation scenario.
@@
Multi-calibration boosting is conducted using the hyperparameter settings listed in Table 2b.
@@ Results
We show the bias of the estimated average treatment effect (ATE) by shift intensity (column panels) and training set size (row panels) for each CATE estimation method in Figure 5 (see also Table 3). The results show that in the present setting all methods are able to produce unbiased estimates of the ATE in the non-shifted test data (first column). Introducing an external shift (second and third column), however, incurs bias across all methods with the shift-reweighted causal forest and shift-reweighted T-learner performing best. The ridge regression-based multi-accurate DR- and T-learner perform best among the shift-blind methods that had no access to the shifted test distribution.
@@
The results for bias of the ATE estimation in Figure 7 (and Table 5) show that in absence of external shift (first column), causal forest-based estimators perform best and are able to achieve unbiasedness. Introducing an external shift between the observational training and test data (second and third columns) amplifies bias such that only the shift-reweighted causal forest is able to approximate the true ATE on average, given sufficient training data. The ridge regression-based multi-accurate DR-learner improve over the initial DR-learner and are competitive with shift-reweighted causal forest in settings with small initial training data and strong shift. Again note that, in contrast to the shift-reweighted methods, the multi-accurate learner had no access to the shifted test distribution during model training.
@@
We plot the bias of the estimated ATE for each method by shift intensity (column panels) and training set size (row panels) in Figure 9 (see also Table 7). In the absence of covariate shift (first column), naive learning in the observational data results in biased estimates of the ATE. Utilizing both data sources in combination via multi-calibration boosting allows to improve over the initial DR- and T-learner. Introducing a covariate shift between the observational data and the RCT (second column) degenerates the performance of the RCT-based estimators and the best results are achieved by multi-accurate DR- and T-learner, especially for strong shifts (third column).
@@
The bias of the estimated ATE for each method by shift intensity (column panels) and training set size (row panels) is presented in Figure 11 (and Table 9). The first set of results (first column) indicate that under unobserved confounding in the observational data only and without external covariate shift, RCT-based estimators are, as expected, unbiased. The multi-accurate DR- and T-learner that draw on both data sources are able to reduce the bias of the naive DR- and T-learner. As the external shift between the observational data and the RCT increases (second and third column), learning only on the RCT incurs bias and shift-reweighted methods as well as (tree-based) multi-accurate DR- and T-learner achieve the best results.
@@
Train Shift CForest S-learner DR-learner T-learner
size degree OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 0.04 0 -0.02 -0.02 0.76 0.35 0.34 -0.09 -0.12 -0.15 -0.11
0.25 1.62 -0.11 2.13 1.87 0.83 1.19 1.39 1.17 0.61 0.64 0.84
0.5 3.17 -0.09 4.2 3.39 1.92 1.88 1.88 2.34 1.34 1.57 1.8
0.75 4.46 -0.55 5.96 4.69 2.43 2.65 3.4 3.41 2.17 2.49 2.45
500 1 5.26 1.52 7.05 5.6 3.69 3.28 3.55 4.24 2.77 3.01 3.45
1.25 5.87 -2.09 7.65 5.92 3.89 3.36 4.42 4.4 2.55 2.77 3.08
1.5 6.29 -2.33 8.1 6.38 2.59 3.68 3.93 5.09 2.46 3.4 3.69
1.75 6.8 -2.73 8.45 7.42 4.28 4.91 4.93 5.13 2.26 3.01 4.21
2 6.88 -4.08 8.6 7.1 5.01 4.01 4.31 5.17 1.69 3.48 3.58
0 0.04 0.04 0.01 0 0.31 -0.01 0.21 -0.01 -0.01 0.16 0.2
0.25 0.79 -0.34 1.79 1.35 0.46 0.61 0.6 0.74 0.33 0.5 0.58
0.5 1.8 -0.65 3.62 2.42 1.41 1.21 1.35 1.68 0.98 1.19 1.24
0.75 2.52 -1.41 5.14 2.87 1.8 1.4 1.92 2.39 1.43 1.77 2
2000 1 2.97 -2.08 6 3.27 1.82 1.51 1.91 2.76 1.58 -2.41 2.18
1.25 3.35 -2.12 6.65 3.96 2.35 1.92 2.1 3.26 1.99 2.2 2.52
1.5 3.63 -2.84 7.07 4.21 2.38 2.19 2.17 3.49 1.92 2.56 2.89
1.75 3.81 -3.27 7.29 4.5 2.29 2.08 2.27 3.59 1.86 2.54 2.84
2 3.93 -3.59 7.48 4.55 2.67 2.32 2.52 3.85 1.71 3.01 3.59
0 0.03 0.02 -0.05 -0.05 0.28 0.14 0.14 -0.07 -0.07 -0.18 -0.09
0.25 0.62 -0.27 1.65 1.15 0.39 0.53 0.43 0.66 0.3 0.48 0.57
0.5 1.31 -0.57 3.41 2.11 0.94 0.99 0.82 1.51 0.84 1.03 1.37
0.75 2.08 -1.22 4.77 2.57 1.5 1.13 1.37 2.25 1.38 1.62 1.81
3500 1 2.38 -1.63 5.64 2.83 1.88 1.54 1.62 2.61 1.55 2 2.29
1.25 2.59 -2.1 6.19 3.03 1.7 1.51 1.87 2.91 1.64 2.16 2.38
1.5 2.72 -3.61 6.45 3.28 2.13 1.59 1.98 3.05 1.61 2.19 2.62
1.75 2.94 -3.18 6.7 3.67 2.18 1.69 1.85 3.23 1.67 2.45 2.73
2 3 -3.97 6.87 3.79 1.9 2.28 2.26 3.35 1.49 2.4 2.87
0 -0.06 -0.05 -0.11 -0.11 -0.06 0.02 -0.01 -0.13 -0.13 -0.14 -0.14
0.25 0.52 -0.25 1.57 1.07 0.32 0.36 0.41 0.62 0.27 0.49 0.71
0.5 1.1 -0.54 3.2 1.82 0.76 0.67 0.7 1.39 0.74 0.96 1.14
0.75 1.65 -1.38 4.56 2.28 1.11 1.11 1.39 2.07 1.2 1.56 1.66
5000 1 2.02 -1.86 5.38 2.55 1.27 1.4 1.59 2.47 1.5 1.83 2.24
1.25 2.34 -1.66 5.91 2.97 1.58 1.69 1.47 2.76 1.69 2.02 2.4
1.5 2.37 -3.08 6.2 2.94 1.93 1.97 1.54 2.89 1.62 2.3 2.23
1.75 2.55 -3.24 6.39 3.26 1.91 1.74 1.45 3.06 1.68 2.49 3.14
2 2.63 -3.56 6.53 3.7 1.83 1.9 1.98 3.2 1.7 2.38 2.62
@@
Train Shift CForest S-learner DR-learner T-learner
size degree OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 12.52 12.5 19.64 19.69 39.11 35.83 34.71 14.99 15.01 23.65 23.34
0.25 16.22 14.87 25.4 24.31 39.71 38.62 33.19 18.39 17.96 13.76 23.99
0.5 23.97 17.31 38.24 30.95 49.65 53.33 34.47 22.98 20.9 15.71 28.8
0.75 33.76 22.6 53.54 39.98 56.69 43.87 48.8 29.32 25.36 21.3 31.88
500 1 39.1 33.58 63.77 46.94 58.16 59.52 49.9 34.82 30.83 22.51 36.65
1.25 44.13 50.24 69.13 47.81 57.9 58.83 60.97 34.24 26.37 22.24 32.55
1.5 47.91 55.52 74.24 51.85 82.21 66.16 56.16 40.65 25.54 24.22 37.26
1.75 53.07 164 78.5 63.76 60.46 64.53 62.66 39.19 29.42 27.65 41.72
2 54.01 100.17 80.29 58.83 68.18 63.73 56.24 40.18 24.46 22.7 37.35
0 5.06 5.1 14.97 14.92 15.75 15.83 15.31 9.56 9.52 8.5 15.34
0.25 6.13 5.98 19.28 17.02 16.96 15.4 15.32 10.43 10.35 8.83 15.77
0.5 10.47 9.33 29.76 20.21 20.28 18.6 17.6 14.83 14.49 11.65 18.63
0.75 13.19 14.76 40.88 21.47 19.86 18.81 19.15 17.24 15.84 13.61 22.23
2000 1 15.44 23.4 47.64 23.43 23.09 21.72 25.05 19.44 17.07 541.18 22.88
1.25 17.03 30.8 53.67 27.69 25.46 22.39 21.78 21.68 19.04 14.56 25.03
1.5 18.42 40.07 57.9 28.71 26.72 26.79 24.99 23.04 17.78 16.42 25.81
1.75 19.23 42.16 59.87 29.85 26.8 22.7 22.47 23.04 16.93 15.57 27.11
2 19.99 45.5 62.18 30.1 24.72 27.07 23.7 25.36 16.63 19.02 29.57
0 3.2 3.19 12.91 12.87 12.56 11.02 11.25 7.64 7.63 7.24 12.44
0.25 4.22 4.17 16.63 14.13 12.04 11.28 10.98 9 9.05 7.34 13.86
0.5 6.57 6.46 26.72 16.98 15.43 13.4 11.72 12.47 12.03 9.51 16.99
0.75 9.78 12.25 36.13 18.59 14.68 14.92 14.56 15.35 14.25 11.21 17.79
3500 1 10.65 14.68 42.75 19.42 17.16 15.14 15.14 16.8 15 12.84 21.89
1.25 10.83 23.62 47.2 19.39 16.18 16.57 16.01 17.78 14.77 12.98 20.17
1.5 11.61 32.26 49.25 20.51 18.26 16.88 17.81 18.68 14.49 13.39 23.67
1.75 12.69 38.13 51.6 22.77 17.92 16.71 16.32 19.75 15.28 15.07 21.18
2 12.73 56.72 53.34 23.36 16.4 21.69 19.04 20.17 15.18 14.76 22.09
0 2.62 2.63 12.45 12.36 9.31 8.73 9.23 6.97 6.99 6.14 10.51
0.25 3.19 3.01 15.53 13.18 9.32 8.49 8.95 7.77 7.73 7.74 12.25
0.5 4.95 4.58 23.96 14.14 10.56 9.48 10.36 10.57 10.09 8.2 14.69
0.75 6.66 9.45 33.15 16.08 12.62 11.07 12.24 13.68 12.92 11.27 16.6
5000 1 7.89 11.84 38.96 16.78 13.9 12.46 12.89 14.89 13.63 11.27 18.55
1.25 9.41 16.77 43.57 18.71 15.24 12.35 13.52 16.02 13.83 22.07 20.03
1.5 9.06 23.63 45.78 17.95 14.93 13.22 13.17 16.68 13.61 13.18 19.91
1.75 9.82 29.24 47.46 19.6 14.82 13.09 14.68 17.55 13.89 14.38 22.76
2 10.36 40.95 48.85 22.73 15.09 14.9 15.18 18.62 14.32 14.42 19.83

### v2

@@ 7 Conclusion
In this work, we connect multi-accurate learning and conditional average treatment effect (CATE) estimation and show how off-the-shelf multi-accurate learning can be used for CATE estimation that is robust to unknown covariate shift. Although we empirically compare to more “state of the art” causal machine learning, these methods were designed for different purposes. Important directions for future work include “best-of-both-worlds” guarantees on both robustness and efficiency by improving variance reduction properties of Multi-CATE. A finer-grained analysis of the statistical implications of algorithmic implementations of boosting could also be relevant, in addition to improving hyperparameter tuning in the causal setting. In our work, we focus on establishing robustness properties.
@@ Appendix A Notation summary
$\mathcal{D}_{\text{obs}}$ Observational dataset
$\mathcal{D}_{\text{rct}}$ Randomized controlled trial dataset
@@ Appendix B Details on algorithms
elsereturn $\tilde{p}=p_{k}$ , rescaling the outcomes // return when miscalibration small
@@ Appendix C Proofs
(2) Suppose any of $\hat{\mu}_{1},\hat{\mu}_{0}$ are not consistent estimators and $e_{1}^{-1},(1-e_{1}(X))^{-1}\in\mathcal{H}.$ The implications of multi-accuracy with respect to $\mathcal{H}$ relate to the classical doubly-robust estimator:
@@ D.1 Data and Software
In all experiments, (initial) T-learner and DR-learner are post-processed using the MCBoost algorithm as implemented in the mcboost (0.4.2) [Pfisterer et al., 2021] package. More concretely, we make use of boosting for degree-2 multi-calibration, a (slightly) stronger notion than multi-accuracy, but computationally less demanding than full multi-calibration Gopalan et al. [2022b]. The hyperparameter settings used for post-processing are listed as part of the following detailed presentation of the experiments (Table 2b and 13b).
@@ D.2.1 Setup
    and used to simulate externally shifted observational data $\mathcal{D}_{os-shift}$ or shifted randomized control trial (RCT) data, $\mathcal{D}_{rct}$ (where $e_{1}(X)=0.5$), depending on the simulation scenario.
@@
To quantify the external shift, we compute the multivariate Kullback-Leibler (KL) divergence [Kullback and Leibler, 1951] between the covariate distribution ($x_{1}\dots x_{10}$) of the training set (simulation 1a, 1b)/ the RCT (simulation 2a, 2b) ($\mathcal{N}^{1}_{p}$) and the test set ($\mathcal{N}^{0}_{p}$).
$\mathrm{I}_{KL}(\mathcal{N}^{0}_{p}\|\mathcal{N}^{1}_{p})=\frac{1}{2}\left\{\mathrm{tr}(\mathbf{\Omega}_{1}\mathbf{\Sigma}_{0})+(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{0})^{\mathrm{T}}\mathbf{\Omega}_{1}(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{0})-p-\ln|\mathbf{\Sigma}_{0}|+\ln|\mathbf{\Sigma}_{1}|\right\},$
where ${\mathbf{\Omega}=\mathbf{\Sigma}^{-1}}$.
@@
Multi-calibration boosting is conducted using the hyperparameter settings listed in Table 2b. Hyperparameter settings of the baseline CATE learner are shown in Table 3.
@@
Table 3: Hyperparameter settings of (baseline) CATE learners. Default settings are used for parameters not listed.
Method Implementation Hyperparameter Value
CForest grf num.trees 2000
mtry sqrt(p)+20
sample.fraction 0.5
honesty.fraction 0.5
min.node.size 5
S-,T-, ranger num.trees 500
DR-learner mtry sqrt(p)
min.node.size 5
replace TRUE
sample.fraction 1
@@ Results
We show the bias of the estimated average treatment effect (ATE) by shift intensity (column panels) and training set size (row panels) for each CATE estimation method in Figure 5 (see also Table 4 which includes quantifications of the distribution shifts using KL divergence). The results show that in the present setting all methods are able to produce unbiased estimates of the ATE in the non-shifted test data (first column). Introducing an external shift (second and third column), however, incurs bias across all methods with the shift-reweighted causal forest and shift-reweighted T-learner performing best. The ridge regression-based multi-accurate DR- and T-learner perform best among the shift-blind methods that had no access to the shifted test distribution.
@@
The results for bias of the ATE estimation in Figure 7 (and Table 6 including KL divergence) show that in absence of external shift (first column), causal forest-based estimators perform best and are able to achieve unbiasedness. Introducing an external shift between the observational training and test data (second and third columns) amplifies bias such that only the shift-reweighted causal forest is able to approximate the true ATE on average, given sufficient training data. The ridge regression-based multi-accurate DR-learner improve over the initial DR-learner and are competitive with shift-reweighted causal forest in settings with small initial training data and strong shift. Again note that, in contrast to the shift-reweighted methods, the multi-accurate learner had no access to the shifted test distribution during model training.
@@
We plot the bias of the estimated ATE for each method by shift intensity (column panels) and training set size (row panels) in Figure 9 (see also Table 8 including KL divergence). In the absence of covariate shift (first column), naive learning in the observational data results in biased estimates of the ATE. Utilizing both data sources in combination via multi-calibration boosting allows to improve over the initial DR- and T-learner. Introducing a covariate shift between the observational data and the RCT (second column) degenerates the performance of the RCT-based estimators and the best results are achieved by multi-accurate DR- and T-learner, especially for strong shifts (third column).
@@
The bias of the estimated ATE for each method by shift intensity (column panels) and training set size (row panels) is presented in Figure 11 (and Table 10 including KL divergence). The first set of results (first column) indicate that under unobserved confounding in the observational data only and without external covariate shift, RCT-based estimators are, as expected, unbiased. The multi-accurate DR- and T-learner that draw on both data sources are able to reduce the bias of the naive DR- and T-learner. As the external shift between the observational data and the RCT increases (second and third column), learning only on the RCT incurs bias and shift-reweighted methods as well as (tree-based) multi-accurate DR- and T-learner achieve the best results.
@@
Train Shift KL- CForest S-learner DR-learner T-learner
size degree Div. OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 0.08 0.04 0 -0.02 -0.02 0.76 0.35 0.34 -0.09 -0.12 -0.15 -0.11
0.25 0.23 1.62 -0.11 2.13 1.87 0.83 1.19 1.39 1.17 0.61 0.64 0.84
0.5 0.65 3.17 -0.09 4.2 3.39 1.92 1.88 1.88 2.34 1.34 1.57 1.8
0.75 1.15 4.46 -0.55 5.96 4.69 2.43 2.65 3.4 3.41 2.17 2.49 2.45
500 1 1.62 5.26 1.52 7.05 5.6 3.69 3.28 3.55 4.24 2.77 3.01 3.45
1.25 1.91 5.87 -2.09 7.65 5.92 3.89 3.36 4.42 4.4 2.55 2.77 3.08
1.5 2.09 6.29 -2.33 8.1 6.38 2.59 3.68 3.93 5.09 2.46 3.4 3.69
1.75 2.22 6.8 -2.73 8.45 7.42 4.28 4.91 4.93 5.13 2.26 3.01 4.21
2 2.40 6.88 -4.08 8.6 7.1 5.01 4.01 4.31 5.17 1.69 3.48 3.58
0 0.02 0.04 0.04 0.01 0 0.31 -0.01 0.21 -0.01 -0.01 0.16 0.2
0.25 0.17 0.79 -0.34 1.79 1.35 0.46 0.61 0.6 0.74 0.33 0.5 0.58
0.5 0.60 1.8 -0.65 3.62 2.42 1.41 1.21 1.35 1.68 0.98 1.19 1.24
0.75 1.11 2.52 -1.41 5.14 2.87 1.8 1.4 1.92 2.39 1.43 1.77 2
2000 1 1.52 2.97 -2.08 6 3.27 1.82 1.51 1.91 2.76 1.58 -2.41 2.18
1.25 1.86 3.35 -2.12 6.65 3.96 2.35 1.92 2.1 3.26 1.99 2.2 2.52
1.5 2.05 3.63 -2.84 7.07 4.21 2.38 2.19 2.17 3.49 1.92 2.56 2.89
1.75 2.21 3.81 -3.27 7.29 4.5 2.29 2.08 2.27 3.59 1.86 2.54 2.84
2 2.32 3.93 -3.59 7.48 4.55 2.67 2.32 2.52 3.85 1.71 3.01 3.59
0 0.02 0.03 0.02 -0.05 -0.05 0.28 0.14 0.14 -0.07 -0.07 -0.18 -0.09
0.25 0.16 0.62 -0.27 1.65 1.15 0.39 0.53 0.43 0.66 0.3 0.48 0.57
0.5 0.60 1.31 -0.57 3.41 2.11 0.94 0.99 0.82 1.51 0.84 1.03 1.37
0.75 1.10 2.08 -1.22 4.77 2.57 1.5 1.13 1.37 2.25 1.38 1.62 1.81
3500 1 1.52 2.38 -1.63 5.64 2.83 1.88 1.54 1.62 2.61 1.55 2 2.29
1.25 1.82 2.59 -2.1 6.19 3.03 1.7 1.51 1.87 2.91 1.64 2.16 2.38
1.5 2.05 2.72 -3.61 6.45 3.28 2.13 1.59 1.98 3.05 1.61 2.19 2.62
1.75 2.20 2.94 -3.18 6.7 3.67 2.18 1.69 1.85 3.23 1.67 2.45 2.73
2 2.29 3 -3.97 6.87 3.79 1.9 2.28 2.26 3.35 1.49 2.4 2.87
0 0.01 -0.06 -0.05 -0.11 -0.11 -0.06 0.02 -0.01 -0.13 -0.13 -0.14 -0.14
0.25 0.16 0.52 -0.25 1.57 1.07 0.32 0.36 0.41 0.62 0.27 0.49 0.71
0.5 0.59 1.1 -0.54 3.2 1.82 0.76 0.67 0.7 1.39 0.74 0.96 1.14
0.75 1.10 1.65 -1.38 4.56 2.28 1.11 1.11 1.39 2.07 1.2 1.56 1.66
5000 1 1.50 2.02 -1.86 5.38 2.55 1.27 1.4 1.59 2.47 1.5 1.83 2.24
1.25 1.81 2.34 -1.66 5.91 2.97 1.58 1.69 1.47 2.76 1.69 2.02 2.4
1.5 2.05 2.37 -3.08 6.2 2.94 1.93 1.97 1.54 2.89 1.62 2.3 2.23
1.75 2.21 2.55 -3.24 6.39 3.26 1.91 1.74 1.45 3.06 1.68 2.49 3.14
2 2.30 2.63 -3.56 6.53 3.7 1.83 1.9 1.98 3.2 1.7 2.38 2.62
@@
Train Shift KL CForest S-learner DR-learner T-learner
size degree Div. OS wOS OS wOS OS Ridge Tree OS wOS Ridge Tree
0 0.08 12.52 12.5 19.64 19.69 39.11 35.83 34.71 14.99 15.01 23.65 23.34
0.25 0.23 16.22 14.87 25.4 24.31 39.71 38.62 33.19 18.39 17.96 13.76 23.99
0.5 0.65 23.97 17.31 38.24 30.95 49.65 53.33 34.47 22.98 20.9 15.71 28.8
0.75 1.15 33.76 22.6 53.54 39.98 56.69 43.87 48.8 29.32 25.36 21.3 31.88
500 1 1.62 39.1 33.58 63.77 46.94 58.16 59.52 49.9 34.82 30.83 22.51 36.65
1.25 1.91 44.13 50.24 69.13 47.81 57.9 58.83 60.97 34.24 26.37 22.24 32.55
1.5 2.09 47.91 55.52 74.24 51.85 82.21 66.16 56.16 40.65 25.54 24.22 37.26
1.75 2.22 53.07 164 78.5 63.76 60.46 64.53 62.66 39.19 29.42 27.65 41.72
2 2.40 54.01 100.17 80.29 58.83 68.18 63.73 56.24 40.18 24.46 22.7 37.35
0 0.02 5.06 5.1 14.97 14.92 15.75 15.83 15.31 9.56 9.52 8.5 15.34
0.25 0.17 6.13 5.98 19.28 17.02 16.96 15.4 15.32 10.43 10.35 8.83 15.77
0.5 0.60 10.47 9.33 29.76 20.21 20.28 18.6 17.6 14.83 14.49 11.65 18.63
0.75 1.11 13.19 14.76 40.88 21.47 19.86 18.81 19.15 17.24 15.84 13.61 22.23
2000 1 1.52 15.44 23.4 47.64 23.43 23.09 21.72 25.05 19.44 17.07 541.18 22.88
1.25 1.86 17.03 30.8 53.67 27.69 25.46 22.39 21.78 21.68 19.04 14.56 25.03
1.5 2.05 18.42 40.07 57.9 28.71 26.72 26.79 24.99 23.04 17.78 16.42 25.81
1.75 2.21 19.23 42.16 59.87 29.85 26.8 22.7 22.47 23.04 16.93 15.57 27.11
2 2.32 19.99 45.5 62.18 30.1 24.72 27.07 23.7 25.36 16.63 19.02 29.57
0 0.02 3.2 3.19 12.91 12.87 12.56 11.02 11.25 7.64 7.63 7.24 12.44
0.25 0.16 4.22 4.17 16.63 14.13 12.04 11.28 10.98 9 9.05 7.34 13.86
0.5 0.60 6.57 6.46 26.72 16.98 15.43 13.4 11.72 12.47 12.03 9.51 16.99
0.75 1.10 9.78 12.25 36.13 18.59 14.68 14.92 14.56 15.35 14.25 11.21 17.79
3500 1 1.52 10.65 14.68 42.75 19.42 17.16 15.14 15.14 16.8 15 12.84 21.89
1.25 1.82 10.83 23.62 47.2 19.39 16.18 16.57 16.01 17.78 14.77 12.98 20.17
1.5 2.05 11.61 32.26 49.25 20.51 18.26 16.88 17.81 18.68 14.49 13.39 23.67
1.75 2.20 12.69 38.13 51.6 22.77 17.92 16.71 16.32 19.75 15.28 15.07 21.18
2 2.29 12.73 56.72 53.34 23.36 16.4 21.69 19.04 20.17 15.18 14.76 22.09
0 0.01 2.62 2.63 12.45 12.36 9.31 8.73 9.23 6.97 6.99 6.14 10.51
0.25 0.16 3.19 3.01 15.53 13.18 9.32 8.49 8.95 7.77 7.73 7.74 12.25
0.5 0.59 4.95 4.58 23.96 14.14 10.56 9.48 10.36 10.57 10.09 8.2 14.69
0.75 1.10 6.66 9.45 33.15 16.08 12.62 11.07 12.24 13.68 12.92 11.27 16.6
5000 1 1.50 7.89 11.84 38.96 16.78 13.9 12.46 12.89 14.89 13.63 11.27 18.55
1.25 1.81 9.41 16.77 43.57 18.71 15.24 12.35 13.52 16.02 13.83 22.07 20.03
1.5 2.05 9.06 23.63 45.78 17.95 14.93 13.22 13.17 16.68 13.61 13.18 19.91
1.75 2.21 9.82 29.24 47.46 19.6 14.82 13.09 14.68 17.55 13.89 14.38 22.76
2 2.30 10.36 40.95 48.85 22.73 15.09 14.9 15.18 18.62 14.32 14.42 19.83